ganeshie8 (ganeshie8):

find \(A^5\) and \(A^{T}\) \(A = B\left(B^{T}B\right)^{-1}B^{T}\) \(B\) is a matrix of size \(m \times n\)

OpenStudy (kainui):

What can you deduce about A from this information alone? How many rows and columns does A have?

ganeshie8 (ganeshie8):

\(\large B^T B\) has n rows and n columns

ganeshie8 (ganeshie8):

\(A\) = (mxn) (nxn) (nxm) so \(A\) is mxm matrix i think

OpenStudy (kainui):

Yeah, so far so good. That tells us a little bit. I guess it makes sense that A is a square matrix since you can't square a matrix unless it's square really.

ganeshie8 (ganeshie8):

oh yeah A^5 requires A to be a square matrx so A^5 and A^T are also a mxm matrices

OpenStudy (kainui):

It might help to think of: \[\LARGE B^TB=C\]

OpenStudy (kainui):

I just mean as a whole matrix to itself. Maybe that's not useful, but apparently in the equation of A they essentially have the inverse of C, which means that B^TB represents an invertible matrix, which is also useful to know.

ganeshie8 (ganeshie8):

yes i believe B^TB must have an inverse eventhough they didn't specify explicitly or is a transpose times a matrix is always invertible ?

OpenStudy (kainui):

\[\LARGE A=B(B^TB)^{-1}B^T\] Well if \[\LARGE \det(B^TB)=0\] Then that means \[\LARGE (B^TB)^{-1}\] is meaningless in the equation for A, right?

OpenStudy (kainui):

Actually it looks like if we left multiply this equation by B^T or right multiply this equation by B we get an interesting result as well!

ganeshie8 (ganeshie8):

\[\LARGE AB=B(B^TB)^{-1}B^TB \] like this ? i was never comfortable with transposes :o

ganeshie8 (ganeshie8):

Oh wait, do you mean to say the last two terms eat each other out ?

OpenStudy (kainui):

Yes exactly!

ganeshie8 (ganeshie8):

\[\LARGE AB=B\color{red}{(B^TB)^{-1}B^TB}\]

ganeshie8 (ganeshie8):

that goes away ? xD

OpenStudy (kainui):

Yeah, so there we have it apparently.

ganeshie8 (ganeshie8):

does that also make A an identity matrix ?

OpenStudy (kainui):

An alternate route is just to look at A^2 \[\LARGE A^2=[B(B^TB)^{-1}B^T][B(B^TB)^{-1}B^T]\]Notice how you can remove the identity matrix because of the inverse and you get: \[\LARGE A^2=B(B^TB)^{-1}B^T = A\] Yes, I believe this means A must be the identity matrix.

OpenStudy (anonymous):

I think it's more easy tan that: \(B^TB=BB^T\) also \(BB^T\) is symetric invertibe matrix that gives the result

ganeshie8 (ganeshie8):

wow! so A^5 = A = I

OpenStudy (kainui):

@myko You can't say this\[\LARGE B^TB = BB^T\] because B is a mxn matrix.

ganeshie8 (ganeshie8):

omg! yes they give different size matrices

OpenStudy (kainui):

In fact, to say that A is the identity matrix is not entirely true. It either has m or n columns, so which is it? =)

ganeshie8 (ganeshie8):

this turned out to be simpler+cuter than it appeared xD thanks a lot !!

OpenStudy (kainui):

Hahaha yes. XD

ganeshie8 (ganeshie8):

oh right, so A is an identity matrix of size m :)

ganeshie8 (ganeshie8):

just a quick question \(A^2 = A \) does not mean \(A = I\) right ? we figured out A is identity matrix from our first observation, by multiplying both sides by B... is this correct ?

OpenStudy (kainui):

Well I think it's a little of both. Ever since that question the other day I've been reviewing everything in linear algebra because my eyes were opened a little. I think there are matrices that aren't the identity that are their own square. And I think there might also be matrices that you can multiply another matrix by to get the same matrix back that isn't the identity. I'm not entirely sure though since I am not sure anymore about anything I can't directly prove yet... So I'll be figuring these out now!

ganeshie8 (ganeshie8):

that makes sense :) atleast we can be sure that \(AB = B\) forces A be an identity

ganeshie8 (ganeshie8):

i think that works irrespective of A or B being invertible..

OpenStudy (akashdeepdeb):

This is from Machine Learning right? This is the algorithm to linear regression!

ganeshie8 (ganeshie8):

i thinkso.. best fit lines right ? it looks like the same matrix : http://en.wikipedia.org/wiki/Idempotent_matrix

ganeshie8 (ganeshie8):

ive just started linear algebra so im bit clueless how that works in linear regression applications :o

ganeshie8 (ganeshie8):

i think that works perfectly if B is a square matrix @OOOPS , then B inverse exists and A equals identity directly

OpenStudy (ikram002p):

i love this question xD good trick

OpenStudy (yanasidlinskiy):

Looks interesting too:D

ganeshie8 (ganeshie8):

\[\large \large AB=B\color{red}{(B^TB)^{-1}B^TB}\] \(\large A\) becomes an Identity only when the matrix we multiply both sides is a multiple of \(B\), in other cases it could be any square matrix. In general, Kainui's second method seems flawless : \[\large A^2 = A = B\left(B^{T}B\right)^{-1}B^{T} = A^{T} \]

OpenStudy (kainui):

Well isn't that matrix you have in red in your last post the identity matrix? Afterall, it is just the product of a matrix and its inverse, it must be the identity matrix right?

ganeshie8 (ganeshie8):

yes, i think we can conlclude A is identity cuz AB=B. and it became identity only because we have multiplied B both sides A need not be an identity if we multiply any other matrix... \[\large \large AX=B\color{red}{(B^TB)^{-1}B^TX}\] not sure if we can make below statement : \[\large AX = X \iff X=cB\]

OpenStudy (ikram002p):

why u study linear algebra anyway ?

OpenStudy (ikram002p):

i think ill review it as well

OpenStudy (ikram002p):

i have a book , and its good also my old notes mmm i used to be good , i think i well do it in reguler since i have no time to see more nline lectures :o

ganeshie8 (ganeshie8):

i am doing lecture 17 as of now and im completely in love with gilbert strang xD they're good for reviewing aswell... 30 lectures will take 1 month @ 1 lecture per day :)

OpenStudy (ikram002p):

haha for u xD for me its hard sometimes to follow english lecture xD lets see if i could match topics that u reach

OpenStudy (ikram002p):

i would give u a copy from my book :-\ if i have soft copy of itt xD even though its the best in linear algebra